OpenStudy (anonymous):

Write the standard equation of a circle with each given radius and center r = 4; C = (-2, 8)

OpenStudy (anonymous):

are u there?

OpenStudy (anonymous):

yes

OpenStudy (shaik0124):

(x-(-2))^2+y-8)^2=4^2

OpenStudy (anonymous):



OpenStudy (anonymous):

so applying the equation of a circle

OpenStudy (shaik0124):

(x+2)^2+(y-8)^2=16

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

i want you to visualize what is going on

OpenStudy (anonymous):

\[(x-(-2))^2+(y-8)^2 = 4^2\]

OpenStudy (anonymous):

the trick with circles is to do with signs

OpenStudy (anonymous):

thank you i understand it now

OpenStudy (anonymous):

\[(x+2)^2+(y-8)^2=16\]

OpenStudy (anonymous):

ok all the best